OpenStudy (anonymous):

find f ' (x) and f '(c) f(x)= 5x^-2 (x+3) value of c = 1

hartnn (hartnn):

is your f(x) =\(5x^{-2}(x+3)\) ??

OpenStudy (anonymous):

yep!

hartnn (hartnn):

distribute 5x^-2 what is 5x^-2 * x = ?

OpenStudy (anonymous):

is this us9ig the constant multiple?

OpenStudy (anonymous):

*using

OpenStudy (anonymous):

5x^-3?

hartnn (hartnn):

x^m.x^n =x^{m+n} so x^-2 . x^1 = x^{-2+1} = ?

OpenStudy (anonymous):

oh darn it's ^-1

hartnn (hartnn):

yes, so u get 5x^-1 and 5*x^(-2)*3 = 15 x^(-2) now do u know the differentiation of x^n ?

OpenStudy (anonymous):

ok is this what i was supposed to do? distribute the (x+3)? if so i get..

hartnn (hartnn):

no? \(\huge\frac{d}{dx}x^n=nx^{n-1}\)

OpenStudy (anonymous):

5x^-1 + 15x^-6 = 20x^-7

hartnn (hartnn):

huh? no! that is not correct. you cannot add exponents when you add terms, you add exponents, only when you multiply terms

OpenStudy (anonymous):

ok, but was 5x^-1 + 15x^-6 correct?

hartnn (hartnn):

and you have f(x) = 5x^(-1) +15x^(-2)

hartnn (hartnn):

how -6 ?

OpenStudy (anonymous):

oh darn. i made the same mistake again for the exponents.

hartnn (hartnn):

f(x) = 5x^(-1) +15x^(-2) now can u differentiate this using the formula i gave ?

OpenStudy (anonymous):

so for 5x^-2 (3) part, i was only supposed to multiply the and 3? and leave the exponent alone? gotcha

OpenStudy (anonymous):

*the 5 and 3

hartnn (hartnn):

yes, because those are constants, x is variable, u leave x alone while multiplying constants

OpenStudy (anonymous):

ok so using that formula i get this:

OpenStudy (anonymous):

-5x^-2 + (-2)15x^-3

hartnn (hartnn):

that is very correct! just need simplification. (-2)*15 = ?

OpenStudy (anonymous):

-30. so then i add both parts together?

hartnn (hartnn):

u get -5x^-2 -30x^-3, right ? now u cannot simplify this further because the exponents of x are different.

OpenStudy (anonymous):

yes, i got those.

OpenStudy (anonymous):

so that's it?

hartnn (hartnn):

yes, that should be it but you can also take -5x^(-2) common.

hartnn (hartnn):

sorry, -5x^(-3) common to get -5x^(-3)(x+6)

OpenStudy (anonymous):

where did the (x+6) come from?

hartnn (hartnn):

-5x^-2 -30x^-3 =-5x^(-3) (x) - 6*5 x^(-3) = -5x^(-3) (x+6)

OpenStudy (anonymous):

ok so i get how you got -5x^-2 -30x^-3

OpenStudy (anonymous):

when I look at =-5x^(-3) (x) - 6*5 x^(-3) i get confused

hartnn (hartnn):

keep it as -5x^-2 -30x^-3 then . now to find f'(c) u just put x=c=1

hartnn (hartnn):

ok?

OpenStudy (anonymous):

so for f'c i get the same thing? since it's jsut 1

hartnn (hartnn):

f'(x)= -5x^-2 -30x^-3 ,right ? we have c=1. so to get f'(1), u just put x=1 in f'(x) that is , put x= 1 in -5x^-2 -30x^-3 make sense ?

OpenStudy (anonymous):

in the back of my book it says: f '(x)= 5[x^-2 (1) + (x+3)(-2x^-3)] =-5(x+6) / x^3

OpenStudy (anonymous):

so that's what you were trying to tell me i guess

hartnn (hartnn):

yes.

hartnn (hartnn):

and u still need tp put x=1

OpenStudy (anonymous):

=-35

hartnn (hartnn):

that is correct :)

OpenStudy (anonymous):

:D